OpenStudy (anonymous):

Ellen paid 11.80 for 40 postage stamps. If some were 33 cents and some were 23 cents how many stamps were 33 cents stamps did she buy ?

OpenStudy (anonymous):

.23x +.33y =11.80 x+y=40 Use substitution: x=40-y .23(40-y) +.33y = 11.80 9.2 - .23y + .33y = 11.80 9.2 + .1y = 11.80 .1y = 2.6 y = 26 Therefore, there are 26 stamps that were 33 cents.